ttttrrrrttttrrr:

Paul deposited $6,460 into a savings account 17 years ago. The account has an interest rate of 4.3% and the balance is currently $13,400.91. How often does the interest compound?

lowkey:

hey, welcome to questioncove (: do you know how to start this

ttttrrrrttttrrr:

just give me the answer

lowkey:

ight, give me money . im helping u, don't be all bossy.

ttttrrrrttttrrr:

lol my bad, you got snapchat

lowkey:

now, let's help you (:

ttttrrrrttttrrr:

sorry ive been stuck on this question for hours im so mad

lowkey:

alright so you need to know this equation \[a=p(1+\frac{ r }{ n })^{nt}\] where a is the amount after the compound or whateber, p is the principal (which is like the starting value), r is the rate of interest, n is the number of years

lowkey:

and t is in time of years

lowkey:

so knowing that info, we can fill in some numbers into our formula

ttttrrrrttttrrr:

yes thats one thing i understand!

lowkey:

ok, try to fill in the formula knowing what u know about the problem and the formula

lowkey:

ya good there lmao

ttttrrrrttttrrr:

soo i think it would be 13,400.91= 6460(1+0.043\[13,400.91= 6460(1+0.043/x) (x times 17)

lowkey:

\[13400.91=6460(1+\frac{ 0.043 }{ x })^{17x}\]

lowkey:

That?

ttttrrrrttttrrr:

thats what ive been putting in my caculator and it keeps popping up 0 lol maybe its broke

lowkey:

You have to isolate the x variable, and that's where it gets tricky

ttttrrrrttttrrr:

lol i got 0 again ughh

lowkey:

Alright lemme help you

ttttrrrrttttrrr:

thnx):

lowkey:

The variable x you put into ur calculator is there for other purposes

lowkey:

Don't use that variable

ttttrrrrttttrrr:

so use y?

lowkey:

So we can do it mathematically at first \[13400.91=6460(1+\frac{ 0.043 }{ x })^{17x}\]

ttttrrrrttttrrr:

ok..

lowkey:

So we can isolate the thing by getting rid of 6460 Divide 13400.91 by 6460

ttttrrrrttttrrr:

i got 2.0744

ttttrrrrttttrrr:

so then what

lowkey:

Lemme type it out. sorry

ttttrrrrttttrrr:

its ok thank you

ttttrrrrttttrrr:

?

ttttrrrrttttrrr:

my computer is going to die

lowkey:

what calculator do you have?

ttttrrrrttttrrr:

a casio

ttttrrrrttttrrr:

is there anyway youll give out your snapchat or no

lowkey:

I mean I guess but I can't help you rn...I gotta go, when is this due?

ttttrrrrttttrrr:

my computer is going to die please add me on snap @ travis_pantazis or instagram @ the same thing PLEASE i need help and my computers going to die

lowkey:

When is this due

ttttrrrrttttrrr:

its not due for a couple of days!

lowkey:

Ok, add my snap and I'll be able to help you by tomorrow afternoon